Find the slope and y-intercept of the line in the given equation y=9/10x-1
Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
y= 9/10x-1 | slope m = (9/10) and the y-intercept is Pt(0,-1)